OpenStudy (darkbeautystar):

Someone please help!!! Misha wants to know how many pepperonis to put on his pizza. He was told that he should use 1 pepperoni rounded to the nearest whole pepperoni for each .75 inches of diameter in a pizza. Show an equation that would solve how many pepperonis he would need if his pizza has a 8 inch radius. Please explain answer!!

OpenStudy (darkbeautystar):

@igreen

OpenStudy (darkbeautystar):

@Hero .. would you please help me??

OpenStudy (darkbeautystar):

@Preetha

OpenStudy (darkbeautystar):

@Michele_Laino

OpenStudy (darkbeautystar):

@jigglypuff314

OpenStudy (darkbeautystar):

@Abhisar

OpenStudy (michele_laino):

since the radius of the pizza is 8 inches, then its circumference \(C\) is: \[\Large C = 2\pi r \cong 2 \cdot 3.14 \cdot 8 = ...?\] since we have 1 pepperoni for each 0.75 inches, then the requested number \(n\) of pepperonis, is: \[\Large n = \frac{C}{{0.75}} = \frac{C}{{\frac{3}{4}}} = C \cdot \frac{4}{3} = ...?\]

OpenStudy (darkbeautystar):

I am not sure @Michele_Laino

OpenStudy (darkbeautystar):

But I am only in 7th grade so I need an answer that suits my grade!!